OpenStudy (anonymous):

Factor the trinomial completely. x^9-17x^8-18x^7

OpenStudy (anonymous):

Take out x^7 as it is the common term x^7(x^2-17x-18) = x^7(x^2-18x+x-18) = x^7(x+1)(x-18) So, x=0,-1,18

OpenStudy (anonymous):

so it would be (x+1)(x-18)?

OpenStudy (anonymous):

No it is x^7(x+1)(x-18)

OpenStudy (anonymous):

Never mind got it thank you :)